Find all vertical and horizontal asymtotes and the x-intercepts. Don't have to graph.
P(x)= (x^2+6x=8) / (x^3+x^2-6x)

Answer by stanbon(48558)

Find all vertical and horizontal asymtotes and the x-intercepts.
----------------------------------------
P(x)= (x^2+6x+8) / (x^3+x^2-6x)
Factor where you can to get:
P(x) = [(x+4)(x+2)] / [x(x+3)(x-2)]
-----------------
Vertical Asymptotes when the denominator is zero: x=0, x=-3, x=2
---------
Horizontal Asymptote: y = 0x^3/x^3 = 0/1 = 0
=============
Cheers,
Stan H.